# BGE M3-Embedding: Multi-Lingual, Multi-Functionality, Multi-Granularity Text Embeddings Through Self-Knowledge Distillation

### 05 Feb 2024 v1

@@ 3.3 Self-Knowledge Distillation
The overall training process is a multi-stage workflow (Figure 2). Firstly, the text encoder is pre-trained with the massive unsupervised data, where only the dense retrieval is trained in the basic form of contrastive learning. The self-knowledge distillation is applied to the second stage, where the embedding model is fine-tuned to establish the three retrieval functionalities. Both labeled and synthetic data are used in this stage, where hard negative samples are introduced for each query following the ANCE method Xiong et al. (2020). Detailed processing is presented in Appendix B.1.
@@ 4.1 Multi-Lingual Retrieval
We evaluate the multi-lingual retrieval performance with MIRACL Zhang et al. (2023c), which consists of ad-hoc retrieval tasks in 18 languages. Each task is made up of query and passage presented in the same language. Following the official benchmark, we evaluate our method using Pyserini Lin et al. (2021), and use nDCG@10 as the primary evaluation metric (Recall@100 is also measured and reported in Appendix C). We incorporate the following baselines in our experiment: the lexical retrieval method: BM25 Robertson and Zaragoza (2009); the dense retrieval methods: mDPR333https://huggingface.co/castorini/mdpr-tied-pft-msmarco Zhang et al. (2023b), mContriever444https://huggingface.co/facebook/mcontriever-msmarco Izacard et al. (2022), mE5 Wang et al. (2022) and E5 Wang et al. (2023). We also make a comparison with Text-Embedding-3-Large(abbreviated as OpenAI-3), which was recently released by OpenAI555https://platform.openai.com/docs/guides/embeddings.
@@ A.2 Synthetic Data
The prompt for GPT3.5 is “You are a curious AI assistant, please generate one specific and valuable question based on the following text. The generated question should revolve around the core content of this text, and avoid using pronouns (e.g., ”this”). Note that you should generate only one question, without including additional content:”. The details of generated datasets are shown in Table 8.
@@ B.3 Split-batch Method
Appendix C Additional Results
In this section, we present additional evaluation results on the MIRACL and MKQA benchmarks. As shown in Table 11 and 12, M3-Embedding outperforms all baselines on average.

### 10 Feb 2024 v3

@@ 3.3 Self-Knowledge Distillation
The overall training process is a multi-stage workflow (Figure 2). We use an XLM-RoBERTa Conneau et al. (2020) model pre-trained further through the RetroMAE Xiao et al. (2022) method as the base text encoder. Firstly, the text encoder is pre-trained with the massive unsupervised data, where only the dense retrieval is trained in the basic form of contrastive learning. The self-knowledge distillation is applied to the second stage, where the embedding model is fine-tuned to establish the three retrieval functionalities. Both labeled and synthetic data are used in this stage, where hard negative samples are introduced for each query following the ANCE method Xiong et al. (2020). Detailed processing is presented in Appendix B.1.
@@ 4.1 Multi-Lingual Retrieval
We evaluate the multi-lingual retrieval performance with MIRACL Zhang et al. (2023c), which consists of ad-hoc retrieval tasks in 18 languages. Each task is made up of query and passage presented in the same language. Following the official benchmark, we evaluate our method using Pyserini Lin et al. (2021), and use nDCG@10 as the primary evaluation metric (Recall@100 is also measured and reported in Appendix C.1). We incorporate the following baselines in our experiment: the lexical retrieval method: BM25 Robertson and Zaragoza (2009); the dense retrieval methods: mDPR333https://huggingface.co/castorini/mdpr-tied-pft-msmarco Zhang et al. (2023b), mContriever444https://huggingface.co/facebook/mcontriever-msmarco Izacard et al. (2022), mE5 Wang et al. (2022) and E5 Wang et al. (2023). To make the BM25 and M3 more comparable, in the experiment, we use the same tokenizer as M3 (i.e., the tokenizer of XLM-Roberta) for BM25. Using the same vocabulary from XLM-Roberta can also ensure that both approaches have the same retrieval latency. The results of BM25 with different tokenizers are shown in Appendix C.2. We also make a comparison with Text-Embedding-3-Large(abbreviated as OpenAI-3), which was recently released by OpenAI555https://platform.openai.com/docs/guides/embeddings.
@@ A.2 Synthetic Data
The prompt for GPT3.5 is “You are a curious AI assistant, please generate one specific and valuable question based on the following text. The generated question should revolve around the core content of this text, and avoid using pronouns (e.g., ”this”). Note that you should generate only one question, without including additional content:”. The details of generated dataset are shown in Table 8.
@@ B.3 Split-batch Method
Appendix C More Results
C.1 Additional Resutls
C.2 Different Tokenizer for BM25
We investigate the impact of different tokenizers on the BM25 method, and the results are shown in Table 11. We can observe that:
Using the Analyzer from Lucene131313https://github.com/apache/lucene/tree/main/lucene/analysis/ can significantly enhance the effectiveness of BM25. Lucene analyzer includes multiple steps typically including tokenization, stemming, stopword removal, etc, achieving better results than directly using the tokenzier of xlm-roberta. Additionally, it’s worth noting that the vocabulary size of the tokenizer from xlm-roberta is limited, resulting in fewer unique tokens after encoding documents (for example, on the MLDR dataset, the tokenizer of xlm-roberta produces 1056 unique terms per article, while Lucene’s analyzer generates 1451 unique terms, which is over 37% more and will increase retrieval latency).
M3 outperforms BM25 models using the same tokenizer on all datasets, indicating that the learned weights are significantly better than the weights calculated by BM25.
The sparse retrieval of M3 outperforms BM25 on Miracl and MKQA datasets. In long document retrieval (MLDR), M3’s sparse doesn’t surpass BM25 but achieves competitive performance. This suggests that BM25 remains a highly competitive baseline model. Exploring tokenizers that perform better for sparse representation is a worthwhile topic for future research.
| method | tokenizer | Miracl | MKQA | MLDR |
| BM25 | Analyzer | 38.5 | 40.9 | 64.1 |
| BM25 | xlm-roberta | 31.9 | 39.9 | 53.6 |
| M3.Sparse | xlm-roberta | 53.9 | 45.3 | 62.2 |
| M3.All | xlm-roberta | 70.0 | 75.5 | 65.0 |